A rectangle is to be 2 m longer than it is wide and have an area of 120 m2. Find its dimensions.

Answer by addingup(3677)

The area of the rectangle is:
Area = Length*Width
The problem says that the length is 2m longer than the width:
L = W+2 In the equation above let's substitute Length with this value:
A = (W+2)*W Or:
W*(W+2) = A Now let's plug in the numbers:
W*(W+2) = 120
W^2+2W = 120 Subtract 120 on both sides
W^2+2W-120 = 0 Now factor the left side (FOIL):
(W-10)(W+12)= 0 Divide into two equations:
W-10= 0 or W+12= 0
W= 10 or W= -12 We know that the number we are looking for is positive, so throw away the -12 and let's try the 10 in the equation:
Length = W+2
L = 10+2= 12 Now for the area:
A= L*W; A= (W+2)*W Plug in the numbers:
120= (10+2)*10
120= 12*10
120= 120 We have the correct answer, our rectangle is 10 wide x 12 long